OpenStudy (chillhill):

Solve for x: 4^x - 6(2^x) + 8 = 0

OpenStudy (anonymous):

\[\left( 2^2 \right)^x-6*2^x+8=0\]

OpenStudy (anonymous):

put\[2^x=t\] squaring \[\left( 2^x \right)^2=t^2\] solve the quadratic for t

OpenStudy (anonymous):

i have missed some steps.You solve in detail.